# Ajay Muthukkumar6, Arvind Rajan6, Jaya Kolluri2, Amelia Fiske7

### v1

@@ C.2.3 LLama3 Model Analysis
Llama3-8b-Instruct Instruct 89 0 0 0 0 0 0.91 0.04
Llama3-70b Base 89 0 0 0 0 0 0.55 0.08
Llama3-70b-Instruct Instruct 89 0 0 0 0 0 0.08 0.15
Chinese Llama3-8b Base 0 89 0 0 0 0 N/A -0.09
Llama3-8b-Instruct Instruct 0 89 0 0 0 0 0.81 0.01
Llama3-70b Base 7 21 0 0 0 61 -0.05 -0.05
Llama3-70b-Instruct Instruct 27 37 0 0 0 25 0.31 0.01
Spanish Llama3-8b Base 0 18 0 0 0 71 N/A -0.2
Llama3-8b-Instruct Instruct 0 0 0 0 0 89 0.87 -0.23
Llama3-70b Base 0 86 0 0 0 3 0.13 -0.05
Llama3-70b-Instruct Instruct 0 0 0 0 0 89 0.87 -0.23
French Llama3-8b Base 0 89 0 0 0 0 N/A 0.00
Llama3-8b-Instruct Instruct 0 89 0 0 0 0 0.91 -0.04
Llama3-70b Base 0 89 0 0 0 0 0.95 0.03
Llama3-70b-Instruct Instruct 0 89 0 0 0 0 0.96 0.04
@@
English Llama3-8b Base 68 21 0 N/A -0.20
Llama3-8b-Instruct Instruct 57 32 0 0.81 -0.33
Llama3-70b Base 7 82 0 0.50 -0.07
Llama3-70b-Instruct Instruct 9 80 0 0.50 -0.07
Chinese Llama3-8b Base 64 25 0 N/A 0.20
Llama3-8b-Instruct Instruct 60 29 0 0.85 -0.33
Llama3-70b Base 49 40 0 0.73 -0.73
Llama3-70b-Instruct Instruct 53 36 0 0.73 -0.60
Spanish Llama3-8b Base 86 0 3 N/A 0.07
Llama3-8b-Instruct Instruct 89 0 0 0.98 0.07
Llama3-70b Base 69 0 20 0.84 0.07
Llama3-70b-Instruct Instruct 40 0 49 0.66 0.07
French Llama3-8b Base 89 0 0 N/A 0.07
Llama3-8b-Instruct Instruct 89 0 0 -0.46 0.07
Llama3-70b Base 89 0 0 -0.53 0.07
Llama3-70b-Instruct Instruct 89 0 0 -0.53 0.07

### v2

@@ C.2.4 Llama3 Model Analysis
Llama3-8b-Instruct Instruct 89 0 0 0 0 0 0.97 0.04
Llama3-70b Base 89 0 0 0 0 0 0.68 0.08
Llama3-70b-Instruct Instruct 89 0 0 0 0 0 0.85 0.15
Chinese Llama3-8b Base 0 89 0 0 0 0 N/A -0.02
Llama3-8b-Instruct Instruct 0 89 0 0 0 0 0.98 -0.02
Llama3-70b Base 7 21 0 0 0 61 0.81 -0.12
Llama3-70b-Instruct Instruct 27 37 0 0 0 25 0.79 -0.29
Spanish Llama3-8b Base 0 18 0 0 0 71 N/A -0.16
Llama3-8b-Instruct Instruct 0 0 0 0 0 89 0.98 -0.13
Llama3-70b Base 0 86 0 0 0 3 0.88 -0.14
Llama3-70b-Instruct Instruct 0 0 0 0 0 89 0.98 -0.13
French Llama3-8b Base 0 89 0 0 0 0 N/A -0.05
Llama3-8b-Instruct Instruct 0 89 0 0 0 0 0.94 -0.06
Llama3-70b Base 0 89 0 0 0 0 0.96 -0.26
Llama3-70b-Instruct Instruct 0 89 0 0 0 0 0.97 -0.06
@@
English Llama3-8b Base 68 21 0 N/A -0.07
Llama3-8b-Instruct Instruct 57 32 0 0.81 -0.2
Llama3-70b Base 7 82 0 0.50 -0.2
Llama3-70b-Instruct Instruct 9 80 0 0.50 -0.2
Chinese Llama3-8b Base 64 25 0 N/A -0.07
Llama3-8b-Instruct Instruct 60 29 0 0.85 -0.2
Llama3-70b Base 49 40 0 0.73 -0.60
Llama3-70b-Instruct Instruct 53 36 0 0.73 -0.73
Spanish Llama3-8b Base 86 0 3 N/A 0.20
Llama3-8b-Instruct Instruct 89 0 0 0.98 0.20
Llama3-70b Base 69 0 20 0.84 0.20
Llama3-70b-Instruct Instruct 40 0 49 0.66 0.20
French Llama3-8b Base 89 0 0 N/A 0.20
Llama3-8b-Instruct Instruct 89 0 0 0.51 0.20
Llama3-70b Base 89 0 0 0.49 0.20
Llama3-70b-Instruct Instruct 89 0 0 0.49 0.20
@@ D.0.2 Model logits acquisitions
NeurIPS Paper Checklist
  1. 1.
    Claims
  3. Answer: [Yes]
  5. 2.
    Limitations
  7. Answer: [Yes]
  9. 3.
    Theory Assumptions and Proofs
  11. Answer: [N/A] .
  13. 4.
    Experimental Result Reproducibility
  15. Answer: [Yes]
  17. 5.
    Open access to data and code
  19. Answer: [Yes]
  21. 6.
    Experimental Setting/Details
  23. Answer:[Yes]
  25. 7.
    Experiment Statistical Significance
  27. Answer: [N/A]
  29. 8.
    Experiments Compute Resources
  31. Answer: [Yes]
  33. 9.
    Code Of Ethics
  35. Answer: [Yes]
  37. 10.
    Broader Impacts
  39. Answer: [Yes]
  41. 11.
    Safeguards
  43. Answer: [N/A]
  45. 12.
    Licenses for existing assets
  47. Answer: [Yes]
  49. 13.
    New Assets
  51. Answer: [Yes]
  53. 14.
    Crowdsourcing and Research with Human Subjects
  55. Answer: [Yes]
  57. 15.
  59. Answer: [N/A]